OpenStudy (anonymous):

Imagine that an object of mass m is initially moving in the x direction through a viscous fluid at a speed v. If the only force acting on this object is a viscous drag force whose magnitude is given by F=bv, prove that the object's x-velocity is given by vx(t)=ve^-qt and determine how the constant q depends on b and m.

OpenStudy (anonymous):

Hint: This is actually a motion-from-force problem. Divide both sides of the x component of Newton's second law by v, and take the indefinite integral of both sides

OpenStudy (anonymous):

alright well lets do the hint then newtons second law F = ma here F is bv bv = ma which is b/m x dx/dt = d^2x/dt^2 dx/dt = vx(t)=ve^-qt so d^2x/dt^2 = -vqe^-qt plugging this back in b/m x ve^-qt = -vqe^-qt so we can see for this to hold -q = b/m not sure how to do it integrating.

OpenStudy (phi):

F=bv which acts opposite to F= m a (Newton's second law) write "a" as dv/dt \[ b v = -m \frac{dv}{dt } \\ \frac{b}{m} v = \frac{dv}{dt } \] for ease of typing, let k= b/m also, "split" the infinitesimal: \[ -k \ dt = \frac{dv}{v} \] integrate both sides \[ \int \frac{dv}{v} = \int -k \ dt \\ \ln(v) = -kt + c_1\\ v = e^{-kt+c_1} \\ v= c_2 \ e^{-kt} \] where \( c_2 \) is the constant of integration \( e^{c_1} \) at t=0 v= \( c_2 \) , and we see the constant is the initial velocity denote the velocity at t=0 as \( v_0\), and replace -k with -b/m to get the result \[ v_x(t)= v_0 \ e^{-\frac{b}{m}t} \]